Al saves pennies. He agreed to give 4/9 of his pennies to Bev if she would give 4/9 of what she got from Al to Carl and if Carl in turn would give 4/9 of what he got from Bev to Dani.​ Bev, Carl, and Dani agreed and Dani received 448 pennies. How many pennies did Al have​ initially?

Answer by greenestamps(13200)



Al initially had x pennies.

Bev got (4/9)x pennies from Al.

Carl got (4/9)((4/9)x) pennies from Bev.

Dani got (4/9)((4/9)((4/9)x)) pennies from Carl.

The number of pennies Dani got was 448:

(4/9)((4/9)((4/9)x)) = 448
(64/729)x = 448
x = 448*(729/64) = 7*729 = 5103

ANSWER: Al had 5103 pennies